# Neural ODEs for holographic transport models without translation symmetry

### v1

@@ Appendix B Training scheme and report
Our Neural ODEs are solved using the Tsitouras 5/4 Runge-Kutta method [71].
@@
To learn the metric, the neural network is initialized by sampling from a standard normal distribution $N(0,1)$. The normal distribution $N(0,0.001)$ with a small standard deviation is utilized for learning the mass and hyperparameter tuning.
@@
There are various hyperparameter tuning methods, which consume different computing resources. Here we propose a simple way to quickly find the exponent $b$ in the square of graviton mass. We define the search range as $[0,4]$ with the step size 1. We use the standard cross-validation method with the simple hold-out technique [72]. We separate the dataset with the frequency range $[0.01,\omega_{\max}]$ into two parts. The part with $[0.01,\omega_{\max}-1]$ goes to the training set and the part with $[\omega_{\max}-0.99,\omega_{\max}]$ to the validation set. The rationale for this separation lies in the dependence of the exponent $b$ on the UV physics. We use the optimizers RMSProp and Adam in order. Their learning rates, epoch numbers, and batch sizes are set as (0.001, 0.0001), (10, 10) and (1, 1). We also add the L1 regularization term $L_{1}=0.1\times\left|\theta\right|$ in the loss function. For each $b$ in the search range, we train 5 times and take the average validation error to select the optimal $b$. As long as $\omega_{\max}$ is large enough, the correct $b$ can be found for the current three targets [^6] We remind that our hyperparameter tuning method has its limitations. Considering the true mass square with the UV behavior $m^{2}=az^{b}+\cdots$, we observe that the required $\omega_{\max}$ increase as $a$ decreases. However, even if $b$ cannot be found accurately, the final learned curve of the mass square usually deviates only slightly at UV. In the future, more sophisticated hyperparameter tuning methods can be explored to improve the results, if necessary.. Specifically, $\omega_{\max}=2$ is large enough for the MG and LA models, and $\omega_{\max}=4$ is large enough for the GA model, see Fig. 8.
@@
The training process can be divided into four stages, which are described as follows.
@@
c) In the first stage, we train $T_{1}=120$ times, each time with randomly initialized trainable parameters. In the $i$-th stage with $i>1$, we obtain the initial parameters from the top $T_{i}=T_{i-1}/2$ trained results with the minimum loss in the $\left(i-1\right)$-th stage. We conduct parallel training on multiple computer cores. When more than $T_{i}/2$ tasks are completed, we stop the $i$-th stage and move on to the $\left(i+1\right)$-th stage. This early stopping can effectively prevent the machine from getting stuck in solving Neural ODEs with bad parameters.
After training, we collect information on the best result, including the minimum loss, the MRE of the learned metric, and the MRE of the learned mass square, see Tab. 1.
Figure 8: Root mean square error (RMSE) for evaluating the hyperparameter $b$ in MG (left), LA (middle) and GA (right) models. Each $b$ is trained five times, yielding five nearly coincident RMSE points in most cases. As one can see, the optimal hyperparameters for three models are 1, 2, and 2, respectively.

### v2

@@ Appendix B Training scheme and report
Our Neural ODEs are solved using the Tsitouras 5/4 method Tsitouras , which is an adaptive and explicit Runge-Kutta method.
@@
To learn the metric, the neural network is initialized by sampling from a standard normal distribution $\mathcal{N}(0,1)$. The normal distribution $\mathcal{N}(0,0.001)$ with a small standard deviation is utilized for learning the mass and hyperparameter tuning.
@@
There are various hyperparameter tuning methods, which consume different computing resources. Here we propose a simple way to quickly find the exponent $b$ in the square of graviton mass. We define the search range as $[0,4]$ with the step size 1. We use the standard cross-validation method with the simple hold-out technique Arlot . We separate the dataset with the frequency range $[0.01,\omega_{\max}]$ into two parts. The part with $[0.01,\omega_{\max}-1]$ goes to the training set and the part with $[\omega_{\max}-0.99,\omega_{\max}]$ to the validation set. The rationale for this separation lies in the dependence of the exponent $b$ on the UV physics. We use the optimizers RMSProp and Adam in order. Their learning rates, epoch numbers, and batch sizes are set as (0.001, 0.0001), (10, 10) and (1, 1). We also add the L1 regularization term $L_{1}=0.1\times\left|\theta\right|$ in the loss function. For each $b$ in the search range, we train 5 times and take the average validation error to select the optimal $b$. As long as $\omega_{\max}$ is large enough, the correct $b$ can be found for the current three targets. Specifically, $\omega_{\max}=2$ is large enough for the MG and LA models, and $\omega_{\max}=4$ is large enough for the GA model, see figure 6.
It should be noted that our hyperparameter tuning method has its limitations. Considering the true mass square with the UV behavior $m^{2}=az^{b}+\cdots$, we observe that the required $\omega_{\max}$ increase as $a$ decreases. However, even if $b$ cannot be found accurately, the final learned curve of the mass square usually deviates only slightly at UV. In the future, more sophisticated hyperparameter tuning methods can be explored to improve the results, if necessary.
Figure 6: The root mean square error (RMSE) for evaluating the hyperparameter $b$ in the MG (left), LA (middle) and GA (right) models. Each $b$ is trained five times, yielding five nearly coincident RMSE points in most cases. As one can see, the optimal hyperparameters for three models are 1, 2, and 2, respectively.
@@
The training process can be divided into multiple stages, which are described as follows.
@@
c) In the first stage, we train $N_{1}=120$ times, each time with randomly initialized trainable parameters. In the $i$-th stage with $i>1$, we obtain the initial parameters from the top $N_{i}=N_{i-1}/2$ trained results with the minimum loss in the $\left(i-1\right)$-th stage. We conduct parallel training on multiple computer cores. When more than $N_{i}/2$ tasks are completed, we stop the $i$-th stage and move on to the $\left(i+1\right)$-th stage. This early stopping can effectively prevent the machine from getting stuck in solving Neural ODEs with bad parameters.
After training, we collect information on the best result, including the minimum loss, the MRE of the learned metric, and the MRE of the learned mass square, see table 1. Note that the MRE is calculated by uniformly sampling 100 points within the allowed range of the radial coordinate $z$.
@@
Appendix C At low temperatures
In order to learn the low-temperature physics, we will modify the algorithm as follows.
1) Loss function
We will use a new loss function
$L=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{2}\left[\left|\frac{\eta_{\mathrm{re}}\left(\omega_{i},\theta\right)}{\bar{\eta}_{\mathrm{re}}\left(\omega_{i}\right)}\right|-1-\log\left|\frac{\eta_{\mathrm{re}}\left(\omega_{i},\theta\right)}{\bar{\eta}_{\mathrm{re}}\left(\omega_{i}\right)}\right|\right],$ (A.26)
which is a combination of relative error and logarithmic difference. It has a natural physical origin and will be elaborated in Ran2024 . Here we introduce it just by experience.
2) Ansatz
We will change the previous ansatz into
$f=(1-z)[4\pi T+(1-z)n_{1}(z,\theta_{1})],$ (A.27)
and
$m^{2}=z^{b}e^{n_{2}(z,\theta_{2})}.$ (A.28)
The former inputs the temperature $T$, which reduces the difficulty of learning near-horizon geometry. The latter is beneficial for learning the mass square in the GA model at low temperatures, where $m^{2}/z^{2}$ grows fast near the horizon. Note that for the low-temperature MG and LA models, as well as all the relatively high-temperature models studied in the main text, the ansatz (A.28) remains applicable, and the performance of Neural ODEs will not be significantly affected. The only price of using eq. (A.28) is to assume that the mass square is positive [^11] Although there is no general argument, the mass square in all the models given in Hartnoll1601 is positive, and the absence of instabilities actually requires the positive mass square in MG and GA models Vegh1301 ; Baggioli2014 ..
3) ODE solver
Since the Tsitouras 5/4 is less stable at low temperatures, we change the solver to the Radau IIA5 Hairer1999 , which is a fully implicit Runge-Kutta method and has high numerical stability.
4) Cutoffs
We set the IR and UV cutoffs as $z_{\mathrm{IR}}=0.999$ and $z_{\mathrm{UV}}=0.001$, by which the low-temperature data can still be generated with sufficient accuracy.
5) Initial values
When learning the mass square of MG and LA models, we initialize the neural network with $\mathcal{N}(0,0.001)$. For other learning targets, we use $\mathcal{N}(0,1)$.
In figures 7-10 and table 2, we generate the data and illustrate the performance of Neural ODEs at low temperatures. We choose the parameters ($5.92\alpha=$ 5.9, $\beta^{2}=5.9$, $\gamma_{1}=4.9$, $\gamma_{2}=1$), which correspond to the same temperature $T\simeq 0.004$.
Figure 7: Training data and performance of Neural ODEs for the MG model at low temperatures.
Figure 8: Training data and performance of Neural ODEs for the LA model at low temperatures.
Figure 9: Training data and performance of Neural ODEs for the GA model at low temperatures.
Figure 10: The RMSE for evaluating the hyperparameter $b$ in three models at low temperatures.
Target $f$ of MG $m^{2}$ of MG $f$ of LA $m^{2}$ of LA $f$ of GA $m^{2}$ of GA
Loss $3\times 10^{-7}$ $<1\times 10^{-16}$ $3\times 10^{-10}$ $<1\times 10^{-16}$ $9\times 10^{-8}$ $2\times 10^{-8}$
MRE $1\times 10^{-3}$ $8\times 10^{-10}$ $2\times 10^{-4}$ $2\times 10^{-9}$ $1\times 10^{-3}$ $1\times 10^{-4}$
Table 2: Minimum loss and MRE of six machine learning experiments at low temperatures.
Appendix D Entanglement entropy
At last, we will use the bulk metric learned by Neural ODEs to predict the derivative of EE with respect to the size of entangling region.
Consider the region $A$ to be a spatial strip with infinite length $L$ and width $l$ on the boundary. The EE of $A$ can be calculated by the RT formula Ryu:2006bv
$S=\frac{\mathrm{Area}(\gamma_{A})}{4G_{N}},$ (A.29)
where $\gamma_{A}$ is the minimal surface in the bulk that is homologous to the boundary region $A$. Using the profile equation for $\gamma_{A}$, the EE and the width can be expressed as Bilson:2010ff
[…]
where $z_{\ast}$ is the radial coordinate at the tip of $\gamma_{A}$, and $\delta$ is the UV cutoff that is introduced to avoid the divergence of EE. Comparing the derivatives of above two equations
[…]
where we have set $\delta\rightarrow 0$, one can obtain a simple formula
$\frac{dS(l)}{dl}=\frac{L}{4G_{N}z_{\ast}^{2}}.$ (A.34)
With these preparations in place, let’s substitute the bulk metric (the learned metric or the true value) into eq. (A.31) and get $l(z_{\ast})$ by numerical integration. By the interpolation, we can construct the inverse function $z_{\ast}(l)$. In terms of eq. (A.34) with $z_{\ast}(l)$, the derivative $S^{\prime}(l)$ can be obtained. In figure (11) and figure (12), we plot $S^{\prime}(l)$ for the three models at (relatively) high temperatures and low temperatures. We also calculate the MRE by uniformly sampling 10000 points within the allowed range of $l$, see table (3).
Figure 11: The performance of Neural ODEs in predicting $S^{\prime}(l)$. From left to right, the three panels represent the results for the MG, LA and GA models at high temperatures.
Figure 12: The performance of Neural ODEs in predicting $S^{\prime}(l)$ for three models at low temperatures.
Model MG LA GA
High temperature $1\times 10^{-5}$ $3\times 10^{-6}$ $2\times 10^{-5}$
Low temperature $1\times 10^{-4}$ $1\times 10^{-5}$ $8\times 10^{-5}$
Table 3: The MRE of $S^{\prime}(l)$ for three models at two temperatures.